# Validation of ML-UQ calibration statistics using simulated reference values: a sensitivity analysis

### v1

@@ III.2 Analysis of the generative model
Having characterized the uncertainty and error distributions, one can put the generative model to the test and try to characterize the generative distribution $D$.
Non-normality of $D$.
In the hypothesis that the generative distribution $D$ is standard normal, one should have approximately $\nu_{eff}\simeq 2k$ (Table 2) for those sets with good $IG$ fits, where $\nu_{eff}$ was estimated for the error distributions (Table 3). This is the case for Set 4, and to a lesser extent for Set 7. On the contrary, Sets 5, 8 and 9 have very large shape parameters which contrast strongly with the small values of $\nu_{eff}$. Reasoning on the single calibrated dataset for which $k$ and $\nu_{eff}$ are available, i.e. Set 9, on should conclude on the non-normality of the generative distribution $D$, or at least that the normality of $D$ cannot be considered as a general hypothesis.
Using $Z$ distribution as a proxy for $D.$
Considering the generative model, the distribution of z-scores for a calibrated dataset should be identical to $D$. To check how close the distribution of z-scores is to the errors generative distribution, simulated error distributions have been generated using the actual uncertainties and two generative distributions: a standard Normal distribution $N(0,1)$ and a unit-variance Student’s-t distribution $t_{s}(\nu_{eff,Z})$ with the degrees of freedom estimated for $Z$ (Table 3). The results are shown in Fig. 5.
Figure 5: Recovery of the error distributions with the generative model, using a normal distribution (red line) and unit-variance Student’s-t with effective degrees of freedom for $Z$ (Table 3; blue line). For Set 6, the Student’s-t model cannot be generated because of the infinite variance for $\nu_{eff,Z}<2$ and it was replaced by $\nu_{eff,Z}=2.1$.
The $t_{s}$-based model offers a better fit to the errors histogram for Sets 2, 5, 6, 7, 8 and 9 (considering that Set 9 is affected by a notable bias of the errors). For the other sets, it is on par with the normal model. Note that for Set 6 the value of $\nu_{eff,Z}$ has been fixed to 2.1 to avoid the infinite variance caused by $\nu_{eff,Z}\leq 2$.
It is thus clear that a normal generative distribution $D=N(0,1)$ is not likely to provide simulated error distributions with properties close to the actual errors, except for a few datasets (Sets 1, 3 and 4). A Student’s-t model for the generative distribution $D=t_{s}(\nu_{eff,Z})$, based on the distribution of z-scores might thus offer an interesting alternative to the normal distribution for the generation of simulated error sets, and it might also be useful to test the sensitivity of simulated statistics to $D$, as tested thereafter.
@@ III.3 Sensitivity analysis
In the previous section, we have seen that if there is a best choice for the generative distribution $D$, it is rarely the normal distribution, which leaves the question of its choice. In absence of strong constraints to guide this choice, it is important to assess the sensitivity to $D$ of the simulated reference values for the candidate calibration statistics, and more globally the sensitivity of the validation $\zeta$-scores.
@@ III.3.1 Sensitivity of $\tilde{\vartheta}_{D,ref}$ to the uncertainty distribution
For ZMS [Fig. 6(a)], the simulation recovers correctly the predefined reference value, showing the consistency of the simulation approach for this statistic. For CC [Fig. 6(b)], $\tilde{\vartheta}_{D,ref}$ depends strongly on the dataset, confirming the absence of a common reference value for these datasets. The ENCE [Fig. 6(c)] and ZMSE [Fig. 6(e)] follow a parallel pattern, which is mostly due to the sensitivity of these variables to the dataset size, $M$ (Appendix B). The correction of this trend by multiplying by $M^{1/2}$ [Fig. 6(d,f)] shows that the common pattern disappears and that each uncertainty set leads to its own reference value, even if a few CIs overlap. The ENCE has an outstanding value of $\tilde{\vartheta}_{D,ref}$ for Set 7, which is not the case for ZMSE, and might be the manifestation of an over-sensitivity of the ENCE to features of the uncertainty distribution.
@@ III.3.2 Sensitivity of $\tilde{\vartheta}_{D,ref}$ to $D$
The simulated reference values are generated for Sets 7 and 8 (as representative of the largest sets with markedly different uncertainty distributions) from their actual uncertainties by using a unit-variance Student’s-t generative distribution $D=t_{s}(\nu)$ for a range of degrees of freedom between 3 and 20, covering the range of values observed for $\nu_{eff,Z}$ (Table 3). The sampling size is $N_{MC}=10^{4}$ and $u_{E}$ is sorted and parted into 50 equal-size bins to estimate the ENCE, ZMSE. The results are presented in Fig. 7.
@@
For ZMS [Fig. 7(a)], one expects to observe the conformity of the simulated reference value with the known reference value (i.e. $\tilde{\vartheta}_{\nu,ref}\simeq 1$), independently of the chosen generative distribution. This is the case, albeit with a notable increase of Monte Carlo sampling uncertainty as $v$ decreases. For CC [Fig. 7(b)], the sensibility to $\nu$ depends on the value of the statistic, and it is less marked for low CC values (Set 8) than for larger ones (Set 7). For the ENCE [Fig. 7(c)] and ZMSE [Fig. 7(d)] statistics, there is no ambiguity, as the simulated reference values depend strongly on $\nu$, with a similar behavior for both statistics and datasets.
@@ III.3.3 Sensitivity of $\zeta$-scores to $D$
The results for the ZMS, CC, ENCE and ZMSE statistics and the pertinent simulation scenarios are reported in Figs. 8-9. The intermediate variables for the calculation of these $\zeta$-scores are reported in Appendix A.
@@
The $\zeta$-scores for the ENCE and ZMSE statistics have been estimated for $N=20$ bins and present very similar features (Fig.9): a notable sensitivity to $D$ and a large difference between the Sim and Sim2 protocols (except for Set 3). The latter observation confirms that one should not use the Monte Carlo CIs to estimate the $\zeta$-scores. Focusing one the SimN and SimT results, it appears that they globally agree to reject consistency, except for Set 3 where consistency would be validated by SimT, as observed in Appendix C.
@@ IV Discussion
To circumvent these issues, it is suggested here to use the bootstrapping-based intervals independently of the existence of a reference value, and to characterize the simulated reference by its standard error (standard deviation of the mean), which is much less sensitive to the choice of $D$. Moreover, with a moderate effort on the Monte Carlo sampling, it is possible to ensure that this standard error becomes negligible before the bootstrapping-based uncertainty, which solves the problem of the dependence of the $\zeta$-scores on the simulated reference uncertainty. The impact of this approach has been shown by comparing SimN to SimT and Sim2N to Sim2T (Figs.8-9): in all cases the proposed approach reduces the sensitivity to $D$, except for CC where it is about identical.
@@ V Conclusion
  • The use of simulated reference values for the CC, ENCE and ZMSE statistics is not reliable if $D$ is unknown. It has been shown that the simulated reference values for these statistics are highly sensitive to the choice of generative distribution $D$. If $D$ is unknown, one cannot estimate a reliable reference value for their validation. For CC, checking that it is positive is the best that can be done and there is not much sense in comparing CC values of different datasets (the largest value is not necessarily the better!). A much more powerful and reliable approach is the plotting of RMSE-based confidence curves with bootstrapped CI and simulated reference value. Similarly, the simulated reference values for measures such as the ENCE and ZMSE are very sensitive to the choice of $D$ and therefore cannot generally be used for validation of conditional calibration. When used for comparative studies, they should be limited to datasets of identical sizes and to identical binning schemes. Validation is possible by alternative approaches, such as the “extrapolation to zero bins” method used in Appendix C, or by the estimation of the “fraction of valid bins” proposed by Pernot(Pernot2023d, ). However, the reliability of the latter approach still needs to be demonstrated for a diversity of datasets.

### v2

@@ III.2 Analysis of the generative model
Considering the generative model, the distribution of z-scores for calibrated datasets should be identical to $D$. To check how close the distribution of z-scores is to the errors generative distribution, simulated error distributions have been generated using the actual uncertainties and two generative distributions: a standard Normal distribution $N(0,1)$ and a unit-variance Student’s-t distribution $t_{s}(\nu_{Z})$ with the degrees of freedom reported in Table 2. The results are shown in Fig. 3.
Figure 3: Recovery of the error distributions with the generative model, using a normal distribution (red line) or a unit-variance Student’s distribution $t_{s}(\nu_{Z})$ ($\nu_{Z}$ from Table 2; blue line). For Set 6, the Student’s-t model cannot be generated because of the infinite variance for $\nu_{Z}<2$ and it was replaced by $\nu_{Z}=2.1$.
The $t_{s}$-based model offers a better fit to the errors histogram for all Sets except 3 and 4, for which it is on par with the normal model. Note that for Set 6 the value of $\nu_{Z}$ had to be increased to 2.1 to avoid the infinite variance due to $\nu_{Z}\leq 2$.
It is thus clear that a normal generative distribution $D=N(0,1)$ is not likely to provide simulated error distributions with properties close to the actual errors, except for a few datasets.
@@ III.3 Sensitivity analysis
In the previous section, we have seen that if there is a best choice for the generative distribution $D$, it is rarely the normal distribution. In absence of strong constraints to guide this choice, it is important to assess the sensitivity to $D$ of the simulated reference values for the candidate calibration statistics, and more globally the sensitivity of the validation $\zeta$-scores.
@@ III.3.1 Sensitivity of $\tilde{\vartheta}_{D,ref}$ to the uncertainty distribution
For the ZMS [Fig. 4(a)], the simulation recovers trivially the predefined reference value, as the simulated z-scores are directly samples of the $D$ distribution. For CC [Fig. 4(b)], $\tilde{\vartheta}_{D,ref}$ is seen to depend strongly on the dataset, confirming the absence of a common reference value for these datasets. The ENCE [Fig. 4(c)] and ZMSE [Fig. 4(e)] follow a parallel pattern, which is mostly due to the sensitivity of these variables to the dataset size, $M$ (Appendix B). The correction of this trend by multiplying the statistics by $M^{1/2}$ [Fig. 4(d,f)] shows that the common pattern disappears and that each uncertainty set leads to its own reference value, even if a few CIs overlap. The ENCE has an outstanding value of $\tilde{\vartheta}_{D,ref}$ for Set 7, which is not the case for the ZMSE, and is certainly a symptom of the stratification of the uncertainties resulting from an isotonic regression post-hoc calibration(Busk2023, ; Pernot2023b_arXiv, ; Pernot2023d, ).
@@ III.3.2 Sensitivity of $\tilde{\vartheta}_{D,ref}$ to $D$
The simulated reference values are generated for Sets 7 and 8 (as representative of the largest sets with markedly different uncertainty distributions) from their actual uncertainties by using a unit-variance Student’s-t generative distribution $D=t_{s}(\nu)$ for a range of degrees of freedom between 3 and 20, covering the range of values observed for $\nu_{Z}$ (Table 2). The sampling size is $N_{MC}=10^{4}$ and $u_{E}$ is sorted and parted into 50 equal-size bins to estimate the ENCE, ZMSE. The results are presented in Fig. 5.
@@
For ZMS [Fig. 5(a)], one expects to observe the conformity of the simulated reference value with the known reference value (i.e. $\tilde{\vartheta}_{\nu,ref}\simeq 1$), independently of the chosen generative distribution. This is the case, albeit with a notable increase of Monte Carlo sampling uncertainty as $v$ decreases. For CC [Fig. 5(b)], the sensitivity to $\nu$ depends on the value of the statistic, and it is less marked for low CC values (Set 8) than for larger ones (Set 7). For the ENCE [Fig. 5(c)] and ZMSE [Fig. 5(d)] statistics, there is no ambiguity, as the simulated reference values depend strongly on $\nu$, with a similar behavior for both statistics and datasets.
@@ III.3.3 Sensitivity of $\zeta$-scores to $D$
The results for the ZMS, CC, ENCE and ZMSE statistics and the pertinent simulation scenarios are reported in Figs. 6-7. The intermediate variables for the calculation of these $\zeta$-scores are reported in Appendix A.
@@
The $\zeta$-scores for the ENCE and ZMSE statistics have been estimated for $N=20$ bins and present very similar features (Fig.7): a notable sensitivity to $D$ and a large difference between the Sim and Sim2 protocols (except for Set 3). The latter observation confirms that one should not use the Monte Carlo CIs to estimate the $\zeta$-scores. Focusing on the SimN and SimT results, it appears that they globally agree to reject consistency, except for Set 3 where consistency would be validated by SimT, as observed in Appendix C. However, we have shown above that for Sets 3 and 4 the SimN scenario is the most plausible.
It is somewhat striking that the validation scores of ENCE and ZMSE are so similar, despite notable differences in the reliability of the RCE and ZMS statistics exposed previously(Pernot2024_arXiv, ). In fact, it was shown that both statistics shared a similar sensitivity to the tailedness of error distributions, unlike their sensitivity to the tailedness of uncertainty distributions, where only the RCE could be notably biased. Binning along $u_{E}$ is expected to erase the tailedness problem of (binned) uncertainty distributions, placing both ENCE and ZMSE on an equal footing. This would not be the case for another binning variable.
@@ IV Discussion
To circumvent these issues, it is suggested here to use the bootstrapping-based intervals independently of the existence of a reference value, and to characterize the simulated reference by its standard error (standard deviation of the mean), which is much less sensitive to the choice of $D$. Moreover, with a moderate effort on the Monte Carlo sampling, it is possible to ensure that this standard error becomes negligible before the bootstrapping-based uncertainty, which solves the problem of the dependence of the $\zeta$-scores on the simulated reference uncertainty. The impact of this approach has been shown by comparing SimN to SimT and Sim2N to Sim2T (Figs.6-7): in all cases the proposed approach reduces the sensitivity to $D$, except for CC where it is about identical.
@@ V Conclusion
  • The use of simulated reference values for the CC, ENCE and ZMSE statistics is not reliable if $D$ is unknown. It has been shown that the simulated reference values for these statistics are highly sensitive to the choice of generative distribution $D$. If $D$ is unknown, one cannot estimate a reliable reference value for their validation. For CC, checking that it is positive is the best that can be done and there is not much sense in comparing CC values of different datasets (the largest value is not necessarily the better!). A much more powerful and reliable approach is the plotting of RMSE-based confidence curves with bootstrapped CI and simulated reference value. Similarly, the simulated reference values for measures such as the ENCE and ZMSE are very sensitive to the choice of $D$ and therefore cannot generally be used for validation of conditional calibration. When used for comparative studies, they should be limited to datasets of identical sizes and to identical binning schemes. Validation is possible by alternative approaches, such as the “extrapolation to zero bins” method used in Appendix C, or maybe by the estimation of the “fraction of valid bins” proposed by Pernot(Pernot2023d, ). However, the reliability of the latter approach still needs to be demonstrated for a diversity of datasets.